One person can do a job in 8 hours. A second person can do it in 12 hours. if the first person works 2 hours less than the second, how many hours will it take them working together?
write an algerair solution. let x= length of time first person works and x+2= length of time second works

Answer by fractalier(2101)

Using t for time, we have
t/8 + (t + 2)/12 = 1 (we use 1 for one job)
Now multiply all by 24 (the LCD) and get
3t + 2(t + 2) = 24
5t + 4 = 24
5t = 20
t = 4
The 1st person works 4 hrs and the second works 6.